OpenStudy (johan14th):

Calculus 2 help please. How do know solve an improper integral to find out if it is convergent or divergent? and ho to know?

OpenStudy (johan14th):

\[\int\limits_{1}^{+infinity}\frac{ x^2 }{ \sqrt{2x^4+1} }dx\] does it converge? Thanks

OpenStudy (phi):

if we factor x^4 out of the radical, and simplify we get \[ \frac{1}{\sqrt{2+\frac{1}{x^4} }} \] in the limit that is 1/sqrt(2) = 0.707 so in the limit you are integrating \[ 0.707 \int_{\text{big number}}^\infty \ dx \]

OpenStudy (johan14th):

How would you factor out? Sorry rusty algebra

OpenStudy (holsteremission):

phi's using the fact that for large \(x\), \[\frac{x^2}{\sqrt{2x^4+1}}\approx\frac{x^2}{\sqrt{2x^4}}=\frac{x^2}{\sqrt2x^2}=\frac{1}{\sqrt2}\]